For each of the following, use an appropriate comparison test to determine the convergence or divergence of the series.

α) Σ 2n +17 22 In n +5
Answer #1

a) The given series is 1 (3 + n)n . Now \frac{1}{(3+n)^n}<\frac{1}{n^n}<\frac{1}{n^2} for n1. Since 2 is convergent, by comparison test 1 (3 + n)n is also convergent.

b) The given series is Σ 2η +17 . Now 2n 2n 2n 17 for n > 2

The series  n Inn is divergent by integral test. That is 3 x Inx . Therefore by comparison test Σ 2η +17 is also divergent.